The company purchased (with cash) and used 77000 yards of raw materials at a cost of $12.70 per yard. Its direct laborers finished goods during the period. Bowens standard cost card for its only product is as follows: worked 40,850 hours and were paid a total of $633,175. The company started and completed 62,000 units of Standard Standard Standard or Boursor Rate(1)(2) Direct materials1 yards 13.00 yard 13.00 Quantity Price Cost per per per 1.0 hours 1s.so r15.50 Direct labor Fixed manufacturing overhead Total standard cost per unit 1.0 hours 23.00 hour23.00 $51.50 Required: 1. When recording the raw material purchases (on account a. The Raw Materials inventory will increase (decrease) by how much? b. The Cash will increase (decrease) by how much? c. The materials price variance will be favorable or unfavorable and by how much? 2. When recording the raw materials used in production: a. The Raw Materials inventory will increase (decrease) by how much? b. The Work in Process inventory will increase (decrease) by how much? c. The materials quantity variance will be favorable or unfavorable and by how much? 3. When recording the direct labor costs added to production: a. The Work in Process inventory will increase (decrease) by how much? b. The Cash will increase (decrease) by how much? c. The labor rate and efficiency variances will be favorable or unfavorable and by how much? 4. When applying fixed manufacturing overhead to production: a. The Work in Process inventory will increase (decrease) by how much? b. The fixed overhead budget and volume variances will be favorable or unfavorable and by how much? 5. When transferring costs from Work in Process to Finished Goods, the Finished Goods inventory will increase (decrease) by how much? Indicate the effect of each variance by selecting Favorable, Unfavorable or None) 1a. The raw materials will 1b. The cash wil 1c. |The mateñals price vañance 2a. The raw materials will 2b. The work in process will by The materials quantity 3a. The work in process 3b. The cash will 3c. The labor rate variance is variance is by The labor efmiciency varanco by 4a. The work in process will 4b. The fixed overhead budget The fixed overhead volume by variance is by 5. The finished goods wil
Answer #1

1a. Increase in raw materials = Actual Quantity * Standard Rate = 77,000 yards * $13 = $1,001,000

1b.Decrease in cash = Actual Quantity * Actual Rate = 77,000 * $12.7 = $977,900

1c. Materials Price Variance = $1,001,000 - $977,900 = $23,100 (Favorable)

2a. Decrease in raw materials = Actual Quantity * Standard Rate = 77,000 yards * $13 = $1,001,000 [ Assuming that all units purchased are used in production]

2b. Increase in Work in Process = Standard Quantity * Standard Rate = (62,000 * 1 yard ) * $13 = $806,000

2c. Material Quantity Variance = $806,000 - $1,001,000 = $195,000 (unfavorable)

3a. Work in Process will increase by = Standard hours * Standard Rate = (62,000 *1 hour ) *15.5 = $961,000

3b. The cash will decrease by = $633,175

3c. Labor Rate Variance = Actual Hours * (Standard Rate- Actual Rate) = 40,850 *( 15.5 - (633175/40850) ) = $0 [Neither favorable or unfavorable]

Labor Efficiency Variance = Standard Rate * ( Standard hours - Actual hours) = 15.5 * ( 62,000 - 40,850) = $327,825 [favorable]

4a. The work in process will increase by = Fixed overhead manufacturing rate per hour * Standard hours = $23 * 62,000* 1 = $1,426,000

4b. Fixed overhead budget variance = Actual fixed overhead - Budgeted fixed overheads = [Data not given in the picture]

Fixed overhead volume variance = [Data not given in the picture]

5. Finished goods will increase by = 62,000 units * $51.5 = $3,193,000